OpenStudy (anonymous):

how would i Find the solution to the equation 64^(4 – x) = 4^2x

OpenStudy (e.mccormick):

You want to get the bases of the powers to be the same using exponent rules.

OpenStudy (anonymous):

so in this instance can i put 2^6(4-x)= 2^2(2x)?

OpenStudy (anonymous):

and then just cancel out the bottom basses

OpenStudy (e.mccormick):

what is \(4^3\)? (makes it easier)

OpenStudy (anonymous):

64

OpenStudy (e.mccormick):

So what is \((4^3)^{4 - x}\)'s exponent when you are all done?

OpenStudy (anonymous):

oh, so do i get them both to be 64? and then i should have 64^(4-x)= 4^3(2x)?

OpenStudy (anonymous):

or 4^3(4-x)?=

OpenStudy (e.mccormick):

Not both to 64. Both to a base of 4. AH HA! There you have it... or what looks like it. \(4^{3(4-x)}\) so \(4^{3(4-x)}= 4^{2x}\)

OpenStudy (anonymous):

alright, so then the basses cancel out right, and then i have just \[3(4-x)=2x\]

OpenStudy (anonymous):

and then simplify by doing 12-3x=2x

OpenStudy (e.mccormick):

Yes! \(4^{3(4-x)}= 4^{2x}\) \(\cancel{4}^{3(4-x)}= \cancel{4}^{2x}\) \(3(4-x)= 2x\)

OpenStudy (e.mccormick):

Yah, at that point it should be easy.

OpenStudy (anonymous):

so then 12=5x

OpenStudy (anonymous):

etc

OpenStudy (e.mccormick):

And there you have it!

OpenStudy (anonymous):

alright so in the end i have 12/5 which is 2.4 is that the correct answer?

OpenStudy (anonymous):

thank you so much for the help by the way!

OpenStudy (e.mccormick):

Yah, 12/5. Have fun!